kai6665:

The great wall of china is 20 feet tall and 15 feet wide. what is the volume of a one-mile section of the wall?

surjithayer:

volume=l*w*h l= 1 mile=1*1760 yards=1760*3=5280 ft width w=15 ft height h=20 ft calculate volume